A flashlight is held at the edge of a swimming pool at a height h = 2.5 m such that its beam makes an angle of θ = 38 degrees with respect to the water's surface. The pool is d = 3.75 m deep and the index of refraction for air and water are n1 = 1 and n2 = 1.33, respectively.

What is the horizontal distance, D, from the edge of the pool to the point on the bottom of the pool where the light strikes? Write your answer in m.
1
Expert's answer
2020-07-23T08:46:25-0400


Firs, let's find the distance "CE = BO". From the triangle "\triangle AOB" expressing the side "BO", obtain:


"BO = \dfrac{AB}{\tan\theta} =h\cot\theta = CE"

Now let's find "EF". According to the Snell's law:


"n_1\sin\theta_1 = n_2\sin\theta_2\\
\sin\theta_2 = \dfrac{n_1\sin\theta_1}{n_2}\\
\theta_2 = \arcsin\left(\dfrac{n_1\sin\theta_1}{n_2}\right)"

From the triangle "\triangle ODF":


"EF = OE\tan\theta_2 = d\tan\theta_2 = d\tan\left(\arcsin\left(\dfrac{n_1\sin\theta_1}{n_2}\right) \right)"

Finally:


"D = CE + EF = h\cot\theta + d\tan\left(\arcsin\left(\dfrac{n_1\sin\theta_1}{n_2}\right) \right)"

Obtain the following number:


"D = 2.5\cot38\degree + 3.75\tan\left(\arcsin\left(\dfrac{1\cdot\sin52\degree}{1.33}\right) \right) \approx 5.83m"

Answer. 5.83 m.